OpenStudy (anonymous):

4[6 − (1 + 2x) + 10] = 2(10 − 3x) show step by step

OpenStudy (anonymous):

I always enjoy using PEMDAS because it helps with multi-step equations like this. P (parentheses) *Brackets always go first then parentheses E (exponents) M (multiplication) D (division) A (addition) S (subtraction(

OpenStudy (anonymous):

Oh, well in this situation parentheses since I think you switched them up, lol. Anyways. You CAN start off with multiplication on the right side. 2(10-3x) becomes 20 - 6x. So that simplifies some of it.

OpenStudy (anonymous):

4[6 − (1 + 2x) + 10] Is it just a 1 there? No variable by it at all?

OpenStudy (anonymous):

4[16-(1+2x)]=2(10-3x) 64-4(1+2x)=20-6x 64-4-8x=20-6x 60-20=8x-6x 40=2x X=40/2 X=20